OpenStudy (anonymous):

A basket contains 15 pieces of fruit: 7 apples, 5 oranges, and 3 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple? A. 6/7 B. 7/45 C. 1/6 D. 4/5

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

\[\frac{5}{15}\times \frac{7}{14}\]

OpenStudy (anonymous):

clean up as \[\frac{1}{3}\times \frac{1}{2}\]